# MS-0001-1922.65

### v1

@@
This review paper provides an in-depth overview of the evolution and advancements in Robust Markov Decision Processes (RMDPs), a field of paramount importance for its role in sequential decision-making amidst uncertainty. Fueled by advances in robust optimization theory and the increasing applications of reinforcement learning techniques, RMDPs literature has been enriched extensively. The review focuses on the formulation of RMDPs, particularly ambiguity sets modeling, which is central to hedging uncertainty. The review systematically classifies the extant methodologies for RMDP formulation into three principal categories: parametric, moment-based, and discrepancy-based approaches, and comprehensively dissects them. The review further delves into the rectangular assumption, which is essential for the computational tractability of RMDPs yet noted for its potential to engender overly conservative policy outcomes. The review summarizes three popular rectangular forms and presents new proof attesting to the NP-hardness of non-rectangular RMDPs. Out of traditional RMDPs scope, the review also surveys recent efforts without conventional rectangular assumptions and burgeoning research trends within the RMDP community. These studies foster the development of more flexible and practical modeling frameworks and enhance the adaptability and performance of RMDPs in the face of uncertainty.
@@
Robust Markov Decision Processes, Ambiguity sets modelling, Reinforcement learning, Rectangularity, New fashions
@@ 1 Introduction
Markov decision processes (MDPs) originate from the seminal works of BELLMAN (1957) and Howard (1960) to model sequential decision-making problems under uncertainty. As powerful mathematical tools, MDPs are widely applied in many areas, such as economics, transportation, inventory control, health care, and manufacture (Parkes & Singh 2003, McKenna et al. 2020, Song & Zipkin 1993, Katehakis & Puranam 2012, Fan et al. 2022, Che et al. 2023). Recently, reinforcement learning (RL), which establishes its theoretical foundations on MDPs (Sutton & Barto 2018), has attracted increasing attention and achieved notable success in practical applications. RL stands out for its potential intelligence on decision-making and active learning capabilities, especially when incorporating deep learning (Matsuo et al. 2022, Kosanoglu et al. 2022, Bushaj et al. 2023, Rolf et al. 2023, Zuccotto et al. 2024). The advancement of RL undeniably paves the way for wider applications of MDPs. However, the realm of sequential decision-making, especially in uncertain and dynamic environments, is fraught with complexities. Naive MDPs presuppose a complete grasp of the underlying dynamics, an assumption that often falls short in real-world situations. With deviated estimation or belief affected by uncertainty, decision-makers should anticipate disappointment on average, owing to the optimization-based selection process. This is the so-called the optimizer’s curse (Smith & Winkler 2006). Even worse, this skewed belief in sequential decision-making will persist throughout the entire planning period, continuing to lead to worse outcomes compared to single-stage decision-making problems.
Motivated by advances in robust optimization theory and practical performance (Scarf et al. 1957, Ben-Tal & Nemirovski 2002, Bertsimas & Sim 2004, Goerigk & Schöbel 2016), robust Markov decision processes (RMDPs) have been investigated for decades. Unlike conventional MDPs, RMDPs are specifically designed to deal with situations where only partial information about the true underlying parameters or probability distributions is available, rather than complete knowledge. The decision-maker can construct a collection of possible parameter values or distributions, called an uncertainty set or ambiguity set, based on the belief that the partial information is credible. Subsequently, the decision-maker can determine the optimal solution under the worst-case scenario over the uncertainty or ambiguity set.
@@
The literature on RMDPs has experienced significant growth and diversification in recent years, marked by substantive contributions from diverse academic fields and a renewed interest in data-driven practices. Researchers have explored RMDP modeling from various aspects, including the exploration of diverse ambiguity sets, the development of effective solving and approximation algorithms, as well as the design of robust algorithms for RL. Furthermore, RMDPs have been combined with other techniques from various perspectives and applied in many areas. These techniques include online learning (Croonenborghs et al. 2007, Badrinath & Kalathil 2021, Cowan et al. 2018), safe learning (Hans et al. 2008, Polo & Rebollo 2011, Wachi & Sui 2020), dynamic risk measure (Bäuerle & Rieder 2014, Chow et al. 2015, Yu & Shen 2022, Rockafellar & Uryasev 2000), and regularization (Farahmand 2011, Zhang et al. 2018), etc. Despite the surge in the literature on RMDPs in recent years, a systematic review of RMDPs formulation remains scarce, with only a handful of tutorials or general reviews available (Mannor & Xu 2019, Keith & Ahner 2021, Badings et al. 2023). This absence of a comprehensive review is a significant gap in the field, as it hinders the advancement of RMDP theory and applications. To bridge this gap, we provide an up-to-date review of the formulation of RMDPs/DRMDPs with uncertain transition matrices. This area has received considerable attention and found widespread application within the fields of management science and operations research. By synthesizing the current state of knowledge in this domain, we aim to establish a solid foundation for future research and practical implementations in RMDPs.
@@ 2 Preliminaries
[…]
@@ 3 Parametric RMDPs
where $p^{L}_{sas^{\prime}}$ and $p^{U}_{sas^{\prime}}$ are the predetermined lower bound and upper bound, respectively. The uncertainty set $\mathcal{P}$ in (5) builds a polyhedron. Consequently, the authors develop a policy iteration algorithm that leverages linear programming techniques in each iteration to effectively address the challenges posed by robust formulations. They prove the proposed algorithm can find an $\epsilon$-optimal policy in a finite number of iterations.
In a continuous line, Givan et al. (2000) investigate a special subclass of MDPIPs, termed as bounded-parameter MDPs (BMDPs). A BMDP $M_{\updownarrow}=(T,\mathcal{S},\mathcal{A},R_{\updownarrow},\mathbf{p}_{\updownarrow},\gamma)$ is similar with the general MDP, but $\mathbf{p}_{\updownarrow}$ and $R_{\updownarrow}$ are closed real intervals instead of real values. Compared to previous work under the general setting, the authors consider a particular MDPs family as the order-maximizing MDPs. By utilizing the special structure of this family, they show that the optimal value function is bounded by BMDPs and directly derive the optimal solution without requiring expensive optimization techniques.
@@
This approach offers a more convenient way to construct ambiguity sets, which can easily achieve better generalization performance when the decision-maker lacks information, compared to explicitly defining intervals for each probability. Additionally, the unique parameter $\alpha$ can convey valuable information, such as the degree of robustness or risk-aware confidence level. In Osogami (2012), the author establishes the connection between RMDPs and risk-sensitive MDPs, where the risk confidence level is set to $1-\alpha$. This relationship provides further insights into the role of ambiguity in decision-making under uncertainty.
@@ 4 Moment-Based RMDPs
While parametric RMDPs can be conducted naturally, determining the rational parameter values range in practice can be challenging. Moreover, imposing constraints on each probability can be restrictive, often resulting in overly conservative or unrealistic solutions. It is reasonable to assume that the decision-maker can access historical observations. In this case, some moment information, such as the mean and variance, may be more readily available than credible parameter value ranges, especially in data-driven paradigms.
@@ 5.1 Norm Ambiguity
Norm ambiguity sets, typically constructed on a probability simplex, measure the discrepancy between two distributions using various norms. While this construction appears analogous to (Satia & Lave Jr 1973), it emphasizes the difference between the entire distributions rather than on individual probabilities. Generally, norm am biguity sets can be represented as follows.
@@
$\mathcal{P}_{s}\triangleq\left\{\mathbf{p}_{s}=(\mathbf{p}_{s1},...,\mathbf{p}_{s|\mathcal{A}|})\in\Delta^{|\mathcal{S}|\times|\mathcal{A}|}\left|\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|},\forall a\in\mathcal{A},\textstyle\sum\limits_{a\in\mathcal{A}}\|\hat{\mathbf{p}}_{sa}-\mathbf{p}_{sa}\|\leq\theta_{s},\theta_{s}\geq 0\right\}\right.,$ (16)
@@
RMDPs with norm ambiguity sets can be easily formulated as convex programs, which makes them a preferred choice for large-scale problems to obtain solutions efficiently. Among the various norms, the $L_{1}$-norm and $L_{\infty}$-norm have garnered significant attention in the literature.
@@ A $\phi$-divergence ambiguity set is defined via
Notably, the literature on RMDPs with $\phi$-divergence ambiguity sets pays more attention to sample complexity and efficient algorithm design in RL recently. In these aspects, there are also interesting and impressive results. We briefly review several advanced literature in the following.
@@ 6 Beyond Rectangularity: Coupled Uncertainty Modelling
Equation (29) could be interpreted as the intersection of $\mathcal{P}$ with affine subspace where $P_{\mathcal{S}^{\prime}}^{0}$ equals $P_{\mathcal{S}^{\prime}}$ from geometric perspective. Without rectangular assumption, $\mathcal{P}_{\mathcal{S}^{*}}(P_{\mathcal{S}^{\prime}})$ obviously depends on $P_{\mathcal{S}^{\prime}}$ and $\mathcal{P}^{\prime}$ for any $\mathcal{S}^{*}\subseteq\mathcal{S}$. Hence, the $k$-rectangular ambiguity sets can be defined rigorously with conditional projection.
@@ 7.1 Satisfactory Regime
RS is a novel framework developed in RO and DRO (Long et al. 2023) that transforms the traditional optimization problem of seeking the optimum into attaining an acceptable target specified ahead. Due to the outstanding interpretability and complete theory of RS, Ruan et al. (2023) extend it into the MDPs version in recent. As opposed to traditional RMDPs with hard constraints to limit the distinction among distributions, RS MDPs impose soft constraints for all other transition distributions. Leveraging the theorems from Long et al. (2023), the authors formulate the original problems as a conic program. The decision variables reflect the magnitude of constraint violation incurred by the distance between the ambiguous transition kernel and the empirical one. Besides, they transform the original problems into a min-max form and propose a primal-dual algorithm to solve the min-max problem efficiently.
@@ 7.3 Risk, Regularization, and Robustness
The concept of weaving regularization into dynamic programming also has been thoroughly investigated within the reinforcement learning field. (Kaufman & Schaefer 2013, Neu et al. 2017, Kostrikov et al. 2021, Ho et al. 2021, Kumar et al. 2022). While these regularized variants of value iteration and policy iteration have been developed, Grand-Clément & Petrik (2022) first propose the convex optimization formulation of RMDPs. This innovative approach integrates an entropic regularization of the robust Bellman operator with a change of variables involving exponential and logarithmic functions. Consequently, algorithms designed for solving convex optimization problems can be merged with this innovative formulation, leading to the development of new algorithms for RMDPs. This novel perspective holds the promise of paving new pathways for future research in this field.

### v3

# Wenfan Ou and Sheng Bi

@@
Fueled by both advances in robust optimization theory and applications of reinforcement learning, robust Markov Decision Processes (RMDPs) have gained increasing attention, due to their powerful capability for sequential decision-making under uncertainty. This review provides an in-depth overview of the evolution and advances in RMDPs formulations, particularly in ambiguity modeling, and classifies these methods for representing uncertainty into three principal approaches: parametric, moment-based, and discrepancy-based, elaborating the trade-offs among the alternative representations. Meanwhile, the review delves into the rectangular assumptions, which guarantee the tractability of RMDPs yet are noted for their conservatism. The review summarizes three popular rectangular conditions and develops a new proof to attest to the NP-hardness of non-rectangular RMDPs. Out of the traditional RMDPs scope, recent efforts without conventional rectangular assumptions and new fashions within the RMDPs community are also reviewed. These studies foster the development of more flexible and practical modeling frameworks and enhance the adaptability and performance of RMDPs.
@@
robust Markov Decision Processes, ambiguity sets modeling, reinforcement learning, rectangularity, new fashions
@@ 1 Introduction
Markov decision processes (MDPs) originate from the seminal works of BELLMAN (1957) and Howard (1960) to model sequential decision-making problems under uncertainty. As powerful mathematical tools, MDPs are widely applied in many areas, such as economics, transportation, inventory control, health care, and manufacture (Parkes & Singh 2003, McKenna et al. 2020, Song & Zipkin 1993, Katehakis & Puranam 2012, Fan et al. 2022, Che et al. 2023). Recently, reinforcement learning (RL), which establishes its theoretical foundations on MDPs (Sutton & Barto 2018), has attracted increasing attention and achieved notable success in practical applications. RL stands out for its potential intelligence on decision-making and active learning capabilities, especially when incorporating deep learning (Matsuo et al. 2022, Kosanoglu et al. 2022, Bushaj et al. 2023, Rolf et al. 2023, Zuccotto et al. 2024). Undeniably, the advances in RL pave the way for wider applications of MDPs. However, volatile and dynamic environments always cause poor performance of MDPs for sequential decision-making. Naive MDPs presuppose a complete grasp of the underlying dynamics, an assumption that often falls short in the real world. With deviated estimation or belief affected by uncertainty, decision-makers should anticipate disappointment on average, owing to the optimization-based selection process. This is the so-called the optimizer’s curse (Smith & Winkler 2006). Even worse, this skewed belief in sequential decision-making will persist throughout the entire planning period, persistently leading to worse outcomes compared to single-stage decision-making problems.
Motivated by advances in robust optimization theory and practical performance (Scarf et al. 1957, Ben-Tal & Nemirovski 2002, Bertsimas & Sim 2004, Goerigk & Schöbel 2016), robust Markov decision processes (RMDPs) have been investigated for decades. Unlike conventional MDPs, RMDPs are specifically designed to deal with situations where only partial information about the true underlying parameters or probability distributions is available, rather than complete knowledge.
The decision-maker can construct a collection of possible parameter values or distributions, called an uncertainty set or ambiguity set, based on the belief that the partial information is credible. Subsequently, the decision-maker can determine the optimal solution under the worst-case scenario over the uncertainty or ambiguity set.
@@
Beneficial from substantial results across diverse fields and interests in data-driven practices, the literature on RMDPs has been significantly enriched in recent years. Researchers have explored RMDP modeling from various aspects, including the exploration of diverse ambiguity sets, the development of effective solving and approximation algorithms, as well as the design of robust algorithms for RL. Furthermore, RMDPs have been combined with other techniques from various perspectives and applied in many areas. These techniques include online learning (Croonenborghs et al. 2007, Badrinath & Kalathil 2021, Cowan et al. 2018), safe learning (Hans et al. 2008, Polo & Rebollo 2011, Wachi & Sui 2020), dynamic risk measure (Bäuerle & Rieder 2014, Chow et al. 2015, Yu & Shen 2022, Rockafellar & Uryasev 2000), and regularization (Farahmand 2011, Zhang et al. 2018), etc. Despite the surge in the literature on RMDPs in recent years, a systematic review of RMDPs formulation remains scarce, with only a handful of tutorials or general reviews available (Mannor & Xu 2019, Keith & Ahner 2021, Badings et al. 2023). We believe that the absence of a specialized review hinders the development of RMDP theory and applications. To bridge this gap, we provide an up-to-date review of the RMDPs/DRMDPs formulations with uncertain transitions. This is a common class of RMDPs problems that have been widely considered in research and applications within the management science and operations research community. By surveying impressive results and studies in this domain, we aim to establish a solid foundation for future research and practical implementations in RMDPs.
@@ 2 Preliminaries
A DRMDP can be represented by a tuple $<T,\mathcal{S},\mathcal{A},\mathcal{P},\gamma>$, where the transition probability matrix $\mathbf{p}$ obeys to a distribution $\mu$, denoted by $\mathbf{p}\sim\mu$. $\mu$ is unknown but belongs to a known distribution family $\mathcal{P}$ of probability distributions, called the ambiguity set. Exactly, DRMDPs can be expressed as
@@ 3 Parametric RMDPs
where $p^{L}_{sas^{\prime}}$ and $p^{U}_{sas^{\prime}}$ are the predetermined lower bound and upper bound, respectively. The uncertainty set $\mathcal{P}$ in (5) builds a polyhedron. Consequently, the authors develop a policy iteration algorithm that leverages linear programming techniques in each iteration to address the challenges posed by robust formulations effectively. They prove the proposed algorithm can find an $\epsilon$-optimal policy in a finite number of iterations.
In a continuous line, Givan et al. (2000) investigate a special subclass of MDPIPs, termed as bounded-parameter MDPs (BMDPs). A BMDP $M_{\updownarrow}=(T,\mathcal{S},\mathcal{A},R_{\updownarrow},\mathbf{p}_{\updownarrow},\gamma)$ is similar with the general MDP, but $\mathbf{p}_{\updownarrow}$ and $R_{\updownarrow}$ are closed real intervals instead of real values. Compared to previous work under the general setting, the authors consider a particular MDPs family as the order-maximizing MDPs. By utilizing the special structure of this family, they show that BMDPs bound the optimal value function and directly derive the optimal solution without requiring expensive optimization techniques.
More recently, Delimpaltadakis et al. (2023) also leverage the threshold structure to expand the scope of BMDPs into continuous action spaces. Distinct from simple real-valued intervals, the authors introduce two functions $f^{L},f^{U}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ to represent the lower and upper bounds on the transition probability with respect to transition $(s,a,s^{\prime})$. Due to strong duality and the ordering threshold structure, they successfully convert the original max-min formulation into a maximization problem. Furthermore, they develop an efficient value iteration algorithm based on this reformulation when $\mathcal{A}$ is a polytope
@@
Contrary to explicitly defining intervals for each probability, this approach just needs to determine a unique parameter $\alpha$ such that the approach can more conveniently construct ambiguity sets and achieve better generalization performance when the decision-maker has limited information about transitions. Additionally, $\alpha$ can be interpreted as the degree of robustness or risk-aware confidence level. In this paper, the author establishes the equivalence between RMDPs and risk-sensitive MDPs, where the risk confidence level is set to $1-\alpha$. This relationship provides further insights to understand and construct ambiguity sets.
@@ 4 Moment-Based RMDPs
While the ideas of parametric RMPDs are natural, it is challenging in practice to determine rational parameter values range. Moreover, parametric approaches impose constraints on each probability that may be too restrictive, often resulting in overly conservative or unrealistic solutions. On the other hand, it is reasonable to assume that the decision-maker can access historical observations. It is reasonable to assume that the decision-maker can access historical observations. In this case, some moment information, such as the mean and variance, may be more readily available than credible parameter value ranges, especially in data-driven paradigms.
@@ 5.1 Norm Ambiguity
Norm ambiguity sets, typically constructed on a probability simplex, measure the discrepancy between two distributions using various norms. While this construction appears analogous to (Satia & Lave Jr 1973), it emphasizes the difference between the entire distributions rather than individual probabilities. Generally, norm ambiguity sets can be represented as follows.
@@
$\mathcal{P}_{s}\triangleq\left\{\mathbf{p}_{s}=(\mathbf{p}_{s1},...,\mathbf{p}_{s|\mathcal{A}|})\in\Delta^{|\mathcal{S}|\times|\mathcal{A}|}\left|\begin{array}[]{l}\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|},\quad\forall a\in\mathcal{A},\\ \textstyle\sum\limits_{a\in\mathcal{A}}\|\hat{\mathbf{p}}_{sa}-\mathbf{p}_{sa}\|\leq\theta_{s},\theta_{s}\geq 0\end{array}\right\}\right.,$ (16)
@@
RMDPs with norm ambiguity sets can be easily formulated as convex programs, which makes them a preferred choice for large-scale problems. Among the various norms, the $L_{1}$-norm and $L_{\infty}$-norm have garnered significant attention in the literature.
@@ A $\phi$-divergence ambiguity set is defined via
Notably, the literature on RMDPs with $\phi$-divergence ambiguity sets has recently paid more attention to sample complexity and efficient algorithm design in RL. In these aspects, there are also interesting and impressive results. We briefly review several advanced literature in the following.
@@ 6 Beyond Rectangularity: Coupled Uncertainty Modelling
Equation (29) could be interpreted as the intersection of $\mathcal{P}$ with affine subspace where $P_{\mathcal{S}^{\prime}}^{0}$ equals $P_{\mathcal{S}^{\prime}}$ from geometric perspective. Without rectangular assumption, $\mathcal{P}_{\mathcal{S}^{*}}(P_{\mathcal{S}^{\prime}})$ obviously depends on $P_{\mathcal{S}^{\prime}}$ and $\mathcal{P}^{\prime}$ for any $\mathcal{S}^{*}\subseteq\mathcal{S}$. Hence, the $k$-rectangular ambiguity sets can be defined rigorously with conditional projection..
@@ 7.1 Satisfactory Regime
RS is a novel framework developed in RO and DRO (Long et al. 2023) that transforms the traditional optimization problem of seeking the optimum into attaining an acceptable target specified ahead. Due to the outstanding interpretability and complete theory of RS, Ruan et al. (2023) extend it into the MDPs version in recent. As opposed to traditional RMDPs with hard constraints to limit the distinction among distributions, RS MDPs impose soft constraints for all other transition distributions. Leveraging the theorems from Long et al. (2023), the authors formulate the original problems as a conic program.
The decision variables reflect the magnitude of constraint violation incurred by the distance between the ambiguous transition kernel and the empirical one. Besides, they transform the original problems into a min-max form and propose a primal-dual algorithm to solve the min-max problem efficiently.
@@ 7.3 Risk, Regularization, and Robustness
The concept of weaving regularization into dynamic programming also has been thoroughly investigated within the reinforcement learning field. (Kaufman & Schaefer 2013, Neu et al. 2017, Kostrikov et al. 2021, Ho et al. 2021, Kumar et al. 2022). While these regularized variants of value iteration and policy iteration have been developed, Grand-Clément & Petrik (2022) first propose the convex optimization formulation of RMDPs. This innovative approach integrates an entropic regularization of the robust Bellman operator with a change of variables involving exponential and logarithmic functions. By conducting this convex reformulation, arbitrary algorithms designed for solving convex optimization problems can be incorporated with the reformulated RMDPs. This fresh perspective has paved the way for the development of innovative algorithms tailored to RMDPs.